OpenStudy (me123):

Could someone please show me how to solve the system by elimination method? 5x+3y=-7 7x-2y=13

OpenStudy (anonymous):

Certainly. The elimination method allows you to form a new equation with only one of your variables by adding a multiple of one equation to another. First pick a variable you want to get rid of.

OpenStudy (me123):

Could you show me how this is done step by step?

OpenStudy (anonymous):

Yes, pick a variable you'd like to eliminate (x or y).

OpenStudy (me123):

x

OpenStudy (anonymous):

Ok, so checking the coefficients on x in both equations, we can see that if we multiply the first equation by \(\frac{-7}{5}\) We will have a -7 coefficient on our x term. So rewrite the first equation after multiplying it by \(\frac{-7}{5}\)

OpenStudy (anonymous):

If anything I say doesn't make sense, stop me and let me know.

OpenStudy (me123):

ok

OpenStudy (anonymous):

So what do you get when you multiply the first equation by \(\frac{-7}{5}\).

OpenStudy (me123):

do you make the 55 a fraction then cross multiply

OpenStudy (anonymous):

No, multiply both sides of the whole equation by -7/5

OpenStudy (anonymous):

\[\frac{-7}{5}(5x+3) = \frac{-7}{5}(-7)\] And simplify.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

While you're working it out I'm gonna grab some coffee.

OpenStudy (me123):

35+3 =-49

OpenStudy (anonymous):

You forgot to (a) change the signs (it was a -7) (b) put each of those over 5 (and simplify the 35) (c) keep your x.

OpenStudy (anonymous):

It should be: \[\frac{-7}{5}(5x + 3y) = \frac{-7}{5}(-7)\] \[\frac{-7}{5}(5x) + \frac{-7}{5}(3y) = \frac{49}{5}\] \[-7x - \frac{21}{5}y = \frac{49}{5}\]

OpenStudy (anonymous):

Right?

OpenStudy (me123):

yes

OpenStudy (anonymous):

Ok, now add this equation to the other one. \(7x\ - 2y = 13\) + \(-7x - (21/5)y = 49/5\) ============== For now don't bother to try to combine the fractions and non-fractions, just eliminate the x.

OpenStudy (me123):

ok

OpenStudy (anonymous):

What do you get for the new equation that only has y's?

OpenStudy (me123):

13/7(7x+2y)=13/7(13)

OpenStudy (me123):

is this step right

OpenStudy (anonymous):

No. Just add the two equations from top to bottom: \(7x \ - 2y \) = 13 + \(-7x - (21/5)y = 49/5\) =============== \(0x - 2y - (21/5)y\) = 13 + 49/5

OpenStudy (me123):

oh ok

OpenStudy (anonymous):

Now multiply that whole equation by 5 to get rid of the fraction

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

The equation on the bottom that only has y's that is.

OpenStudy (me123):

so if i multiply all by 5 I get 10y -10=65+245

OpenStudy (anonymous):

No. 5*(-2y) = -10y 5*(-21/5)y = -21y 5*13 = 65 (this part you did correctly) 5*49/5 = 49 So the equation would be -10y -21y = 65 + 49

OpenStudy (anonymous):

Then you combine your terms.

OpenStudy (me123):

-31y=114

OpenStudy (anonymous):

Right, and now y =?

OpenStudy (anonymous):

Just leave it as a fraction.

OpenStudy (me123):

-31/114

OpenStudy (anonymous):

Yep. Now plug that in for y into either one of your original equations and solve for x.

OpenStudy (anonymous):

Err wait, no. It's -114/31

OpenStudy (anonymous):

You need to practice your distribution rules, and working with equations. I suggest http://www.khanacademy.org/video/solving-equations-with-the-distributive-property?playlist=Algebra%20I%20Worked%20Examples

OpenStudy (me123):

thank you

OpenStudy (anonymous):

Of course! =)

OpenStudy (amistre64):

tag! your it